Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments--Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 12,000 $ 16,200 $ 28,200
Estimated variable manufacturing overhead per machine-hour $ 2.20 $ 3.00
Job P Job Q
Direct materials $ 21,000 $ 12,000
Direct labor cost $ 27,400 $ 10,700
Actual machine-hours used:
Molding 2,500 1,600
Fabrication 1,400 1,700
Total 3,900 3,300

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1-8, assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

1. What was the company’s plantwide predetermined overhead rate? (Round your answer to 2 decimal places.)

-How much manufacturing overhead was applied to Job P and how much was applied to Job Q? (Do not round intermediate calculations.)

-What was the total manufacturing cost assigned to Job P? (Do not round intermediate calculations.)

-If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)
Answer #1

1) Plantwide predetermine overhead rate :

Estimated manufacturing overhead = (2500*2.2+1500*3+28200) = $38200

Predetermine overhead rate = 38200/4000 = $9.55 per machine hour

2) Calculate manufacturing overhead applied

Job P Job Q
Overhead applied 3900*9.55 = $37245 3300*9.55 = $31515

3) Calculate total manufacturing cost applied job P :

Job P
Direct material 21000
Direct labor 27400
Overhead applied 37245
Total manufacturing cost 85645

4) Unit product cost = 85645/20 = $4282 per unit